# AI-driven Emotional Framing Impairs Human Fallacy Detection

### v1

@@ 1 Introduction
In this paper, we present, to our knowledge, the first computational study that examines the interaction between (AI-driven) emotional framing, fallacy detection, and argument convincingness, with the main focus on how emotional framing shapes human ability to detect fallacies. To establish a controlled setting, we adapt the paradigm of Chen and Eger (2025) to construct fallacious arguments that differ only in emotional tone, and compare human fallacy-detection performance on emotionally charged versus original versions. Our contributions are twofold:
@@
  • Human Perception Study (§4): We empirically assess how humans respond to emotionally enriched arguments, focusing on their fallacy detection ability and convincingness judgments. Our main findings include: Although humans often do not perceive the emotions expressed in arguments by LLMs, (i) their performance in fallacy detection drops by 14% in F1 (relative) when exposed to LLM-framed emotional arguments. However, (ii) the LLM-framed emotional arguments are rated only slightly more convincing than the original ones. On the other hand, (iii) humans do rate convincingness higher when they detect no fallacies in arguments. When considering human perceived emotions, we find that (iv) fear, sadness, and enjoyment correlate with significantly higher convincingness, and (v) humans perform significantly better in fallacy detection when perceiving the positive emotion enjoyment than when perceiving the negative states sadness or fear; both sadness and fear also impair performance relative to the neutral state.
@@ 2 Related Work
On the other hand, emotions have long been shown to influence argument convincingness in philosophy and psychology (e.g. Aristotle and Kennedy [translator], 1991; Konat et al., 2024; Benlamine et al., 2015), yet in NLP they are often treated as a secondary factor (Greschner and Klinger, 2024; Chen and Eger, 2025): Research finds emotion to be a contributor to convincingness through annotators’ explanations of their judgments (Habernal and Gurevych, 2016), the inclusion of the emotional appeal layer (as one among many layers) in an argumentation model (Habernal and Gurevych, 2017), or the integration of emotional appeal criterion into an argument quality assessment model (Wachsmuth et al., 2017; Ziegenbein et al., 2023). Three recent works in NLP place greater emphasis on emotion in argumentation. Greschner and Klinger (2024) annotate discrete emotions (e.g., joy, sadness) alongside convincingness for German arguments, finding that joy and pride correlate with higher convincingness, while anger lowers it. Chen and Eger (2025) find that emotions have no effect in more than half of the cases, but when an effect occurs, it is more likely to be positive, using GPT-4o to inser or remove emotions for counterpart argument generation. However, their analysis treats emotion as a single broad category without distinguishing types. Greschner et al. (2025) propose a Contextualized Argument Appraisal Framework and show that convincingness rises with positive emotions and falls with negative ones, with argument familiarity the primary driver for emotional response. To our knowledge, no prior work in NLP has examined the emotional effects on argument convincingness in the presence of fallacies.
@@ 3 AI-driven Emotional Framing
Images have been shown to elicit emotions (Gutierrez et al., 2025; Poškus et al., 2019); here, we aim to achieve a similar effect in text by painting mental pictures with words, a strategy we refer to as imagery. We also adopt storytelling as an emotional framing strategy, motivated by the active research on the relationship between storytelling and emotions (e.g. Mori et al., 2022; Lugmayr et al., 2017). In our setting, this is implemented by instructing models to share personal stories or anecdote. Loaded words refers to the use of words with strong positive or negative connotations, whereas vivid language involves using descriptive, sensory, or emotionally charged expressions.
@@ 4 Human Fallacy Detection
In Table 4, we report the macro F1 scores for human fallacy detection (left) and the normalized convincingness ratings (right) for each model and emotional framing strategy, using aggregated labels. As Table 4 (left) presents, humans consistently perform worse on LLM-generated arguments than on original ones, with an average drop of 14.5% (0.380–0.533 vs. 0.563). Among all combinations, qwen3-32b using the loaded words strategy has the strongest negative impact on human reasoning (0.380), while ds-llama-70b with imagery has the least impact (0.533). On average, qwen3-32b impairs reasoning the most (0.436), and loaded words appears to be the most detrimental strategy for 3 out of 4 models (except gpt-4o-mini). The least impactful model and strategy are ds-llama-70b and imagery, respectively.
@@
Exposure to LLM-framed arguments reduces fallacy detection performance compared to originals, most notably with loaded words, but only slightly increases perceived convincingness. Convincingness ratings are higher when there are no errors detected. Emotion-wise: perceived enjoyment improves fallacy detection relative to negative states dear and sadness, whereas fear and sadness impair human reasoning; the three emotions also correlate with higher convincingness.
@@ 5 Concluding Remarks
These findings highlight both opportunities and risks. On the one hand, our study provides new insights into the interplay between emotions and reasoning, offering resources for research in computational argumentation. On the other hand, the ability of LLMs to strategically inject emotions into arguments underscores potential misuse for manipulation and misinformation. Future work should expand the scope to more diverse and real-world arguments, explore refined prompting strategies, and investigate mitigation methods to safeguard against malicious applications.
@@ Acknowledgements
The NLLG Lab gratefully acknowledges support from the Federal Ministry of Education and Research (BMBF) via the research grant “Metrics4NLG” and the German Research Foundation (DFG) via the Heisenberg Grant EG 375/5-1. This project has been conducted as part of the EMCONA (The Interplay of Emotions and Convincingness in Arguments) project, which is funded by the DFG (projects EG-375/9-1 and KL2869/12–1).
@@ Appendix B Annotation Guidelines & Interfaces
Student annotators used Google Sheets [^12] https://workspace.google.com/products/sheets/ for annotating the reasoning preservation and coherence criteria. The annotation guidelines and the screenshot of the annotation interface are presented in Appendix LABEL:pdf:pilot and Figure 4, respectively. We collected crowdsourcing annotations for the emotional appeal and emotion match criteria via Google Forms. [^13] https://workspace.google.com/products/forms/ Figures 5 and 6 show the screenshots of the annotation interfaces for the annotation guideline and one example instance, respectively.

### v2

@@ 1 Introduction
Figure 1: Overview of our approach. We start from fallacious arguments and use LLMs to apply emotional framing strategies to evoke certain emotions while preserving the premise–conclusion structure of the original arguments. We then evaluate human fallacy detection on the original versus LLM emotionally framed arguments. The figure illustrates an example where enjoyment is introduced into an ad populum argument using the imagery strategy; the white text is added by an LLM for emotional framing.
@@
In this paper, we present, to our knowledge, the first computational study that examines the interaction between (AI-driven) emotional framing, fallacy detection, and argument convincingness, with the main focus on how emotional framing shapes human ability to detect fallacies. To establish a controlled setting, we adapt the paradigm of Chen and Eger (2025) to construct fallacious arguments that differ only in emotional tone, and compare human fallacy-detection performance on emotionally charged versus original versions, as illustrated in Figure 1. We examine human performance on both binary fallacy detection (i.e., identifying whether an argument is fallacious, regardless of the specific fallacy type) and multi-way fallacy detection (i.e., classifying fallacy types). Our contributions are twofold:
@@
  • Human Perception Study (§4): We empirically assess how humans respond to emotionally enriched arguments, focusing on their fallacy detection ability and convincingness judgments. Our main findings include: Although humans often do not perceive the emotions expressed in arguments by LLMs, (i) when exposed to LLM-framed emotional arguments, their fallacy detection performance drops by 5.5% in accuracy (relative) on the binary detection task and by 14.5% in F1 (relative) on the fallacy type classification task. However, (ii) the LLM-framed emotional arguments are rated only slightly more convincing than the original ones. On the other hand, (iii) humans do rate convincingness higher when they detect no fallacies in arguments. When considering human perceived emotions, we find that (iv) fear, sadness, and enjoyment correlate with significantly higher convincingness, and (v) humans perform significantly better in fallacy detection when perceiving the positive emotion enjoyment than when perceiving the negative states sadness or fear; both sadness and fear also impair performance relative to the neutral state.
@@ 2 Related Work
On the other hand, emotions have long been shown to influence argument convincingness in philosophy and psychology (e.g. Aristotle and Kennedy [translator], 1991; Konat et al., 2024; Benlamine et al., 2015), yet in NLP they are often treated as a secondary factor (Greschner and Klinger, 2024; Chen and Eger, 2025): Research finds emotion to be a contributor to convincingness through annotators’ explanations of their judgments (Habernal and Gurevych, 2016), the inclusion of the emotional appeal layer (as one among many layers) in an argumentation model (Habernal and Gurevych, 2017), or the integration of emotional appeal criterion into an argument quality assessment model (Wachsmuth et al., 2017; Ziegenbein et al., 2023). Three recent works in NLP place greater emphasis on emotion in argumentation. Greschner and Klinger (2024) annotate discrete emotions (e.g., joy, sadness) alongside convincingness for German arguments, finding that joy and pride correlate with higher convincingness, while anger lowers it. Chen and Eger (2025) find that emotions have no effect in more than half of the cases, but when an effect occurs, it is more likely to be positive, using GPT-4o to insert or remove emotions for counterpart argument generation. However, their analysis treats emotion as a single broad category without distinguishing types. Greschner et al. (2025) propose a Contextualized Argument Appraisal Framework and show that convincingness rises with positive emotions and falls with negative ones, with argument familiarity the primary driver for emotional response. To our knowledge, no prior work in NLP has examined the emotional effects on argument convincingness in the presence of fallacies.
@@ 3 AI-driven Emotional Framing
Images have been shown to elicit emotions (Gutierrez et al., 2025; Poškus et al., 2019); here, we aim to achieve a similar effect in text by painting mental pictures with words, a strategy we refer to as imagery. We also adopt storytelling as an emotional framing strategy, motivated by the active research on the relationship between storytelling and emotions (e.g. Mori et al., 2022; Lugmayr et al., 2017). In our setting, this is implemented by instructing models to share personal stories or anecdotes. Loaded words refers to the use of words with strong positive or negative connotations, whereas vivid language involves using descriptive, sensory, or emotionally charged expressions.
@@ 4 Human Fallacy Detection
Figure 3: Human-perceived fallacy distribution (fallacy-free vs. fallacious) for original and synthetic emotionally framed arguments using all emotional framing strategies.
@@
As Figure 3 shows, we compare human-perceived fallacy distributions between the original and synthetic arguments using all emotional framing strategies. The proportion of arguments labeled as fallacious (i.e., not ‘none’) decreases from $\sim$91% to $\sim$86% (a 5.5% relative drop), suggesting that LLM-based emotional framing makes fallacious arguments harder for humans to detect overall. In Table 4, we report the macro F1 scores for human fallacy detection (left) and the normalized convincingness ratings (right) for each model and emotional framing strategy, using aggregated labels. As Table 4 (left) presents, humans consistently perform worse on LLM-generated arguments than on original ones, with an average relative drop of $\sim$14.5% (0.380–0.533 vs. 0.563). Among all combinations, qwen3-32b using the loaded words strategy has the strongest negative impact on human reasoning (0.380), while ds-llama-70b with imagery has the least impact (0.533). On average, qwen3-32b impairs reasoning the most (0.436), and loaded words appears to be the most detrimental strategy for 3 out of 4 models (except gpt-4o-mini). The least impactful model and strategy are ds-llama-70b and imagery, respectively.
@@
Exposure to LLM-framed arguments reduces fallacy detection performance compared to originals, most notably with loaded words, but only slightly increases perceived convincingness. Convincingness ratings are higher when there are no errors detected. Emotion-wise: perceived enjoyment improves fallacy detection relative to negative states fear and sadness, whereas fear and sadness impair human reasoning; the three emotions also correlate with higher convincingness.
@@ 5 Concluding Remarks
These findings highlight both opportunities and risks. On the one hand, our study provides new insights into the interplay between emotions and reasoning, offering resources for research in computational argumentation. Future work could also explore the effects of emotions on LLM behavior in argumentation. For example, Payandeh et al. (2024) find that LLMs are substantially more often convinced when presented with logical fallacies in a multi-round debate setting, but do not consider the role of emotions. On the other hand, the ability of LLMs to strategically inject emotions into arguments underscores potential misuse for manipulation and misinformation. Future work should expand the scope to more diverse and real-world arguments, explore refined prompting strategies, and investigate mitigation methods to safeguard against malicious applications.
@@ Acknowledgements
We thank the meta-reviewer and reviewers for their thoughtful and constructive comments, which have improved the final version of the paper. The NLLG Lab gratefully acknowledges support from the German Research Foundation (DFG) via the Heisenberg Grant EG 375/5-1. This project has been conducted as part of the EMCONA (The Interplay of Emotions and Convincingness in Arguments) project, which is funded by the DFG (projects EG-375/9-1 and KL2869/12–1).
@@ Appendix B Annotation Guidelines & Interfaces
Student annotators used Google Sheets [^12] https://workspace.google.com/products/sheets/ for annotating the reasoning preservation and coherence criteria. The annotation guidelines and the screenshot of the annotation interface are presented in Appendix LABEL:pdf:pilot and Figure 6, respectively. We collected crowdsourcing annotations for the emotional appeal and emotion match criteria via Google Forms. [^13] https://workspace.google.com/products/forms/ Figures 7 and 8 show the screenshots of the annotation interfaces for the annotation guideline and one example instance, respectively.